# GMIMS-DRAGONS: A Faraday Depth Survey of the Northern Sky Covering 350 to 1030 MHz

### v1

@@ 3.2 Calibration Observations
Intensity calibration proceeded in three stages: (1) we used a calibration noise source to correct gain fluctuations, (2) we observed raster scans of bright radio sources at the beginning and end of each survey night to calibrate the data in Jy beam-1, and (3) we used beamwidth measurements to convert the units from Jy beam-1 into brightness temperature (K). These procedures were similar to those used for the GMIMS-HBN survey (Wolleben et al., 2021). We did not use any polarized on-sky sources for polarization angle calibration, but found good agreement between the DRAGONS polarization angles and previous datasets without the need for angle corrections (Section 5.2).
3.2.1 Injected Noise Signal
@@
The noise-source signal was injected into the feed with an integrated calibration noise coupler (Flygare et al., 2017). The noise source was switched on for 6 seconds every 3 minutes while the telescope was scanning, synchronized with the recorded integrations. We randomized the start time of the first noise-source instance for each scan in order to avoid coherent patterns of flagged data in the maps. We used the recorded noise-source signals to correct outside-temperature-dependent gain variations in the signal chain over the course of each observing night. The azimuth scans as well as the calibration raster scans (see Section 3.2.2) were initially converted into noise-source units, as described in Section 4.2.1. For this purpose, the reliance on the stability of the noise source was minimal, requiring only that its output should be stable through one night. However, for the survey-wide calibrations that were ultimately applied, the noise source was required to be stable over the course of the entire survey, and we found this criterion to be adequately met (Section 4.2.2).
3.2.2 Raster Scans
@@
Figure 11: An example of the raw data in a ${e}_{\text{low}}$ scan from Phase 1 of the survey. (a) The total power, RR* and LL*, and (b) the polarized signal, RL* and LR*, as a function of azimuth at 610 MHz in spectrometer units (binned into a 1 MHz wide channel). (c) The RA and declination coverage as a function of azimuth for this particular ${e}_{\text{low}}$ scan. (d) The total power and (e) polarized signal frequency spectrum in spectrometer units at ${a}=350\arcdeg$. Please note the factor of 10 difference in displayed scale between the total (panels a, d) and polarized (panels b, e) intensities. The regularly spaced, rectangular shaped signals in the azimuth plots are the noise source firing (note this signal exceeds the power scale in panel b). In panel e we show the persistent RFI mask in blue to mitigate confusion between the frequency-dependent fluctuations and the RFI spikes in the polarization data.
As the four calibrators are unresolved in the $1.3\arcdeg$ to $3.6\arcdeg$ beam of the DRAO-15, in the raster scans the flux density of each is convolved across the Gaussian of the beam, resulting in a two-dimensional (2D) Gaussian representing the source. For each raster scan, we used a Levenberg-Marquardt least-squares algorithm, included in the astropy.modeling package (Astropy Collab. et al., 2022), to fit a 2D Gaussian plus a sloped planar background to each frequency channel in both LL* and RR*. Cyg A required a second Gaussian to be added to the model to account for nearby Cyg X in order to accurately characterize the background. An example of the resulting fit for a Cyg A scan at 1010 MHz is shown in Figure 10. We compared the amplitude of the fitted Gaussian for each calibrator in noise-source units to the known flux density and spectral index of the calibrators (Perley & Butler, 2017) to determine the noise source to Jy beam-1 conversion factor (see Section 4.2.2).
In addition to the amplitudes, the outputs from the fitting process included the HPBW of the Gaussian and its position. The HPBW values from such a scan are shown in Figure 2, and are used in the absolute calibration step (Section 4.8) to convert the data into units of brightness temperature (K). The positions obtained served as a daily check of telescope pointing (Figure 4); no significant changes in pointing were detected in the course of the survey.
@@
Reflector optics Shaped offset Gregorian with Mizigutch condition
@@
Spatial sampling Full Nyquist (2807 of 2880 planned scans)
Observation dates 2022 June to 2023 February
Data loss to RFI  $\sim{30}$% of frequency band
@@
3.3 The Observed Raw Dataset
@@
An example of a raw azimuth scan is shown in Figure 11, and we note a few key features here. The noise-source signal appears as regularly spaced, rectangular-shaped pulses in the azimuth plots (panels a and b). There is an offset of $\sim 20\%$ between the RR* and LL* total power (panel a), which is corrected once the calibration to Jy beam-1 is applied to each of RR* and LL* independently (Sections 3.2.24.2.2). Instrumental polarization causes an offset in both the RL and LR products (panel b), which is corrected in the initial calibration stage (Section 4.3). The strong source near ${a}=150\arcdeg$ (panel a) is the Galactic midplane, which produces instrumental leakage in the polarized signal (panel b). We quantify this leakage using the calibrated maps in Section 5.3. The increase in polarized signal between ${a}=345\arcdeg$ and ${a}=15\arcdeg$ (the ends of the scan; panel e) corresponds to the highly polarized Fan Region, centered on Galactic coordinates $\ell\approx 130\arcdeg$, $b\approx 5\arcdeg$ (Hill et al., 2017).
4 Data Processing
In this section we describe the steps in processing the survey data, from applying the calibrations described in Section 3.2 through to the final maps. These steps are outlined in Figure 12. We processed the raw scans, which were stored as hdf5 files, using python for RFI flagging (Section 4.1) and calibration into Jy beam-1 (Section 4.2), after which we converted them to SDFITS files for the map-making steps. The latter included ground (Section 4.4) and ionosphere (Section 4.5) correction, followed by destriping and combining East and West maps (Section 4.7), for which we used fortran and IDL routines developed by Carretti et al. (2019) for the S-band Polarization All Sky Survey (S-PASS). We refer to this set of map-making software as ‘the S-PASS software’.
@@ 4 Data Processing
Figure 12: Flowchart depicting the main data processing steps. Blue boxes (rows 1 and 2) represent the raw data in spectrometer units (Section 3.3). Green boxes (rows 3 and 4) correspond to the calibration steps (Sections 4.2.14.2.2). Yellow (rows 5 – 7) and orange boxes (rows 8 and 9) represent various amplitude (Sections 4.34.4) and angle (Section 4.5) correction steps. Red boxes (rows 10 – 14) show the map-making steps (Section 4.6).
Figure 13: An example of the RFI masking for a ${e}_{\text{low}}$ scan from Phase 1 of the survey. The waterfall plot shows the polarized intensity, PI, converted into units of Jy beam-1. The persistent RFI mask is shown in grey, and the intermittent RFI mask for this scan is shown in black. The plots to the left and at the bottom show one-dimensional cuts indicated by the blue lines on the waterfall plot: the vertical cut shows the frequency spectrum (left) at a particular timestamp, and the horizontal cut shows the time series data (bottom) at 837 MHz
4.1 Radio Frequency Interference
@@
Although the DRAO site is largely protected from RFI by the surrounding mountains as well as government regulation, contamination from wireless communications still affects a significant portion of the DRAGONS frequency band. We applied two stages of RFI excision to the 83 kHz channelized survey data prior to feeding into the S-PASS software: (i) identifying persistent contaminated channels (e.g. TV and mobile phone bands) and (ii) flagging each scan for instances of intermittent RFI. We excluded channels with persistent RFI from use in further processing stages. If intermittent RFI contaminated only a small fraction of the scan, we calibrated the scan and included it in the map-making. The S-PASS map-making software also includes a step for removing any remaining single-pixel RFI instances, as well as discarding scans with high rms variability, which can be indicative of remaining, low-level RFI (Section 4.3).
@@
We identified channels with persistent RFI from the Gaussian fits to the calibration raster scans. For each channel of a raster scan, we determined a $\chi^{2}$ value from the residuals of the fitted model to the scan data. If the $\chi^{2}$ for a channel was consistently high across all phases of the survey, the channel was flagged as containing persistent RFI. This eliminated 25% of the channels, and those remaining were subject to intermittent RFI detection in the individual scans. This persistent RFI mask is shown in grey for one example azimuth scan in Figure 13.
@@
In each azimuth scan we identified intermittent RFI as segments of data characterized by a sharp change in signal level as a function of time or frequency in total or polarized intensity (PI). In each frequency channel in the scan, we identified candidate RFI events by flagging integrations with a PI signal three times the median PI along the scan and grouping together contiguous instances into a candidate event. For each candidate RFI event, we then verified that it was indeed RFI by finding the edge of the event in the frequency domain, scanning within 5 MHz on either side of the central frequency to find where the slope of PI versus frequency exceeded a set threshold. This avoided the misidentification of brightly polarized regions on the sky as RFI. Independently, we checked for large changes in total intensity along the frequency axis for each integration to pick up RFI instances missed by the time-domain scanning approach. Examples of intermittent RFI and the results of the flagging algorithm are shown in black for the example azimuth scan in Figure 13, demonstrating their extent in time and frequency.
@@
From the persistent and intermittent RFI masks, we produced a combined mask for each azimuth scan. We ultimately binned the data into 1 MHz bins for map-making (Section 4.3) and the subsequent Faraday synthesis. Any 83.3 kHz channels identified as containing persistent RFI were excluded from the binning, while channels with only intermittent RFI were included, with the contaminated integrations masked. With twelve 83.3 kHz channels contributing to each 1 MHz bin, it was possible to retain useful data from these narrow channels even in cases where more than 50% of integrations were masked for intermittent RFI, provided the adjacent channels were not heavily contaminated. In some cases, a 1 MHz bin is on the edge of a wider band of persistent RFI, resulting in only a few channels contributing to the bin. In these cases, the 1 MHz channels suffer increased noise (Section 5.3) and a few were discarded following visual inspection of the resulting maps (Section 6.1).
4.2 Calibrating the Scans
@@
4.2.1 Conversion to Noise Source Units
@@
Here $\text{PI}_{\text{noise}}$ is the ‘polarized intensity’ of the noise polynomial, $\text{PI}_{\text{noise}}=\sqrt{(\text{RL*}_{\text{noise}})^{2}+(\text{LR*}_{\text{noise}})^{2}}$ and $\tau_{\text{noise}}=\frac{1}{2}\tan^{-1}(\text{RL*}_{\text{noise}},\text{LR*}_{\text{noise}})$ is the corresponding phase. Following the conversion of the data into noise-source units, we masked the integrations that included the noise-source signal for subsequent processing steps. The Gaussian fits to the raster scans (Section 3.2.2) were then done in noise-source units (Figure 10).
4.2.2 Conversion to Jy beam-1
@@
We determined the noise-source-units to Jy beam-1 conversion factor as a function of frequency by comparing the Gaussian-fitted amplitudes of the raster scans in RR* and LL* to the flux density of the corresponding calibrator sources in Perley & Butler (2017).
@@
Two of the calibrators, Cas A and Tau A, are supernova remnants, and are known to change over time. We tested their suitability as calibrators using the Cyg A conversion factors to calibrate the Cas A and Tau A raster amplitudes for days when both Cyg A and one of these were observed. Cyg A is a bright radio galaxy and its flux is not expected to vary in time. From the resulting calibrated Cas A and Tau A spectra, we found that these sources have changed substantially since the observations for Perley & Butler (2017) and we rejected them for calibration purposes. When the fourth calibrator, the radio galaxy Vir A, was calibrated using the conversion factors developed for Cyg A, the resulting spectrum agreed with Perley & Butler (2017). However, the lower flux density of Vir A led to a noisy spectrum, and using conversion factors developed from Vir A raster scans would have introduced additional noise to the scans. Ultimately, we only used Cyg A raster scans to calibrate all azimuth scans.
@@
We produced survey-wide conversion factors by combining the fitted amplitudes for all Cyg A raster scans to produce a single set of RR* and LL* amplitude spectra. For the conversion spectra to be valid for all scans, the noise source was required to remain constant throughout the survey. Microwave noise sources are sensitive to temperature, and so the noise source was maintained at a controlled temperature above ambient (at $30\arcdeg$ C) in a well-insulated box. We tested the effectiveness of this in achieving the required stability by applying the master conversion factor spectra to all Cyg A and Vir A raster scans and finding the average ratio between the expected spectrum for the source from Perley & Butler (2017) and the measured source spectrum for each scan. Figure 14 shows the average ratio as a function of ambient temperature for each Cyg A (blue dots) and Vir A (red dots) scan. We note that although there is a slight correlation of this ratio with ambient temperature, the ratio remains within 3.5% of unity across the full range of temperatures. To account for the slight variation in the noise source across the survey, we applied a temperature-based scale factor to each scan. We determined the temperature scale factor from the linear fit to the average expected-to-measured ratio as a function of temperature values, shown as a solid black line in Figure 14.
@@
We applied the noise-source-units to Jy beam-1 conversion factors to the RR* and LL* components of all azimuth scans such that each of RR* and LL* represent the total intensity. After this calibration process the median percent difference between RR* and LL* across all scans was less than 5% across most of the frequency band, peaking at 10% near 450 MHz. We then calculated the Stokes $I$ total intensity values for the azimuth scans as $I=\frac{1}{2}(\text{RR*}+\text{LL*})$, following Wolleben et al. (2019, 2021). [^8] The correct definition of Stokes $I$ is $I=\text{RR*}+\text{LL*}$, without the factor of 1/2 (Robishaw & Heiles, 2021). However, our method of calibrating RR* and LL* intensities independently produces maps of total and polarized intensity that agree in brightness with previous radio surveys.
@@
At this stage, no further angle corrections were applied to RL* and LR*, and it was assumed that $Q=\text{RL*}$ and $U=\text{LR*}$ (similar to Wolleben et al., 2019, 2021). The noise-source units-to-Jy beam-1 conversion factor, $R_{QU}$, used for Stokes $Q$ and $U$ is
@@
Figure 14: The ratio of the expected Perley & Butler (2017) flux density to the raster scan amplitudes calibrated using the survey-wide calibration factors averaged across the spectrum for each scan as a function of ambient temperature. Vir A scans are plotted with red squares and Cyg A scans are plotted with blue circles. A linear fit to the ratio versus temperature plot is shown as a black line.
@@
4.3 Base level Adjustments and Filtering
Following RFI excision and calibration into Jy beam-1, we binned the Stokes $Q$, $U$ and $I$ data into 1 MHz frequency channels, each spanning twelve of the 83.3 kHz-wide channels. At this stage, we also separated the scans into East and West halves, which we then saved as SDFITS files to feed into the S-PASS software. Henceforth we use ‘scan’ to refer to an individual $180\arcdeg$ East or West scan rather than the full $360\arcdeg$ raw observation scans. We applied several initial steps to the full set of scans at each binned frequency to clean up the data prior to inclusion in the maps:

### v2

@@ III.2 Calibration Observations
Intensity calibration proceeded in three stages: (1) we used a calibration noise source to correct gain fluctuations, (2) we observed raster scans of bright radio sources at the beginning and end of each survey night to calibrate the data in Jy beam-1, and (3) we used beamwidth measurements to convert the units from Jy beam-1 into brightness temperature (K). These procedures were similar to those used for the GMIMS-HBN survey (Wolleben et al., 2021). We did not use any polarized on-sky sources for polarization angle calibration, but found good agreement between the DRAGONS polarization angles and previous datasets without the need for angle corrections (Section V.2).
III.2.1 Injected Noise Signal
@@
The noise-source signal was injected into the feed with an integrated calibration noise coupler (Flygare et al., 2017). The noise source was switched on for 6 seconds every 3 minutes while the telescope was scanning, synchronized with the recorded integrations. We randomized the start time of the first noise-source instance for each scan in order to avoid coherent patterns of flagged data in the maps. We used the recorded noise-source signals to correct outside-temperature-dependent gain variations in the signal chain over the course of each observing night. The azimuth scans as well as the calibration raster scans (see Section III.2.2) were initially converted into noise-source units, as described in Section IV.2.1. For this purpose, the reliance on the stability of the noise source was minimal, requiring only that its output should be stable through each night. However, for the survey-wide calibrations that were ultimately applied, the noise source was required to be stable over the course of the entire survey, and we found this criterion to be adequately met (Section IV.2.2).
III.2.2 Raster Scans
@@
Reflector optics Shaped offset Gregorian with
Mizigutch condition
@@
Spatial sampling Full Nyquist (2807 of 2880
planned scans)
Observation dates 2022 June to 2023 January
Data loss to RFI  $\sim{25}$% of frequency band
@@
As the four calibrators are unresolved in the $1.3\arcdeg$ to $3.6\arcdeg$ beam of the DRAO-15, in the raster scans the flux density of each is convolved across the Gaussian of the beam, resulting in a 2D Gaussian representing the source. For each raster scan, we used a Levenberg-Marquardt least-squares algorithm, included in the astropy.modeling package (Astropy Collab. et al., 2022), to fit a 2D Gaussian plus a sloped planar background to each frequency channel in both LL* and RR*. Cyg A required a second Gaussian to be added to the model to account for nearby Cyg X in order to accurately characterize the background. An example of the resulting fit for a Cyg A scan at 1010 MHz is shown in Figure 10. We compared the amplitude of the fitted Gaussian for each calibrator in noise-source units to the known flux density and spectral index of the calibrators (Perley and Butler, 2017) to determine the noise-source-to-Jy beam-1 conversion factor (see Section IV.2.2).
In addition to the amplitudes, the outputs from the fitting process included the HPBW of the Gaussian and its position. The HPBW values from such a scan are shown in Figure 2, and are used in the absolute calibration step (Section IV.8) to convert the data into units of brightness temperature (K). The positions obtained served as a daily check of telescope pointing (Figure 4); no significant changes in pointing were detected in the course of the survey.
III.3 The Observed Raw Dataset
@@
Figure 11: An example of the raw data in a ${e}_{\text{low}}$ scan from Phase 1 of the survey. (a) The total power, RR* and LL*, and (b) the polarized signal, RL* and LR*, as a function of azimuth at 610 MHz in spectrometer units (binned into a 1 MHz wide channel). (c) The RA and declination coverage as a function of azimuth for this particular ${e}_{\text{low}}$ scan. (d) The total power and (e) polarized signal frequency spectrum in spectrometer units at ${a}=350\arcdeg$. Please note the factor of 10 difference in displayed scale between the total (panels a, d) and polarized (panels b, e) intensities. The regularly spaced, rectangular-shaped signals in the azimuth plots are the noise-source firing (note this signal exceeds the power scale in panel b). In panel e we show the persistent RFI mask in blue to mitigate confusion between frequency-dependent fluctuations and the RFI spikes in the polarization data.
@@
Figure 12: Flowchart depicting the main data processing steps. Blue boxes (rows 1 and 2) represent the raw data in spectrometer units (Section III.3). Green boxes (rows 3 and 4) correspond to the calibration steps (Sections IV.2.1IV.2.2). Yellow (rows 5 – 7) and orange boxes (rows 8 and 9) represent various amplitude (Sections IV.3IV.4) and angle (Section IV.5) correction steps. Red boxes (rows 10 – 14) show the map-making steps (Section IV.6).
An example of a raw azimuth scan is shown in Figure 11, and we note a few key features here. The noise-source signal appears as regularly spaced, rectangular-shaped pulses in the azimuth plots (panels a and b). There is an offset of $\sim 20\%$ between the RR* and LL* total power (panel a), which is corrected once the calibration to Jy beam-1 is applied to each of RR* and LL* independently (Sections III.2.2IV.2.2). Instrumental polarization causes an offset in both the RL and LR products (panel b), which is corrected in the initial calibration stage (Section IV.3). The strong source near ${a}=150\arcdeg$ (panel a) is the Galactic midplane, which produces instrumental leakage in the polarized signal (panel b). We quantify this leakage using the calibrated maps in Section V.4. The increase in polarized signal between ${a}=345\arcdeg$ and ${a}=15\arcdeg$ (the ends of the scan; panel e) corresponds to the highly polarized Fan Region, centered on Galactic coordinates $\ell\approx 130\arcdeg$, $b\approx 5\arcdeg$ (Hill et al., 2017).
IV Data Processing
In this section we describe the steps in processing the survey data, from applying the calibrations described in Section III.2 through to the final maps. These steps are outlined in Figure 12. We processed the raw scans, which were stored as hdf5 files, using python for RFI flagging (Section IV.1) and calibration into Jy beam-1 (Section IV.2), after which we converted them to SDFITS files for the map-making steps. The latter included ground (Section IV.4) and ionosphere (Section IV.5) correction, followed by destriping and combining east and west maps (Section IV.7), for which we used fortran and IDL routines developed by Carretti et al. (2019) for the S-band Polarization All Sky Survey (S-PASS). We refer to this set of map-making software as ‘the S-PASS software’.
IV.1 Radio Frequency Interference
@@
Although the DRAO site is largely protected from RFI by the surrounding mountains as well as government regulation, contamination from wireless communications still affects a significant portion of the DRAGONS frequency band. We applied two stages of RFI excision to the 83 kHz channelized survey data prior to feeding into the S-PASS software: (i) identifying persistent contaminated channels (e.g. TV and mobile phone bands) and (ii) flagging each scan for instances of intermittent RFI. We excluded channels with persistent RFI from use in further processing stages. If intermittent RFI contaminated only a small fraction of the scan, we calibrated the remaining data in the scan and included the scan in the map-making. The S-PASS map-making software also includes a step for removing any remaining single-pixel RFI instances, as well as discarding scans with high rms variability, which can be indicative of remaining, low-level RFI (Section IV.3).
@@
We identified channels with persistent RFI from the Gaussian fits to the calibration raster scans. For each channel of a raster scan, we determined a $\chi^{2}$ value from the residuals of the fitted model to the scan data. If the $\chi^{2}$ for a channel was consistently high across all phases of the survey, the channel was flagged as containing persistent RFI. This eliminated 25% of the channels, and those remaining were subject to intermittent RFI detection in the individual scans. This persistent RFI mask is shown in gray for one example azimuth scan in Figure 13.
Figure 13: An example of the RFI masking for a ${e}_{\text{low}}$ scan from Phase 1 of the survey. The waterfall plot shows the linearly polarized intensity, PI, converted into units of Jy beam-1. The persistent RFI mask is shown in gray, and the intermittent RFI mask for this scan is shown in black. The six vertical stripes of high PI are the calibration noise source. The plots to the left and at the bottom show one-dimensional cuts indicated by the blue lines on the waterfall plot: the vertical cut shows the frequency spectrum (left) at a particular timestamp, and the horizontal cut shows the time series data (bottom) at 837 MHz
@@
In each azimuth scan we identified intermittent RFI as segments of data characterized by a sharp change in signal level as a function of time or frequency in total or linearly polarized intensity (PI). In each frequency channel in the scan, we identified candidate RFI events by flagging integrations with a PI signal 3 times the median PI along the scan and grouping together contiguous instances into a candidate event. For each candidate RFI event, we then verified that it was indeed RFI by finding the edge of the event in the frequency domain, scanning within 5 MHz on either side of the central frequency to find where the slope of PI versus frequency exceeded a set threshold. This avoided the misidentification of brightly polarized regions on the sky as RFI. Independently, we checked for large changes in total intensity along the frequency axis for each integration to pick up RFI instances missed by the time-domain scanning approach. Examples of intermittent RFI and the results of the flagging algorithm are shown in black for the example azimuth scan in Figure 13, demonstrating their extent in time and frequency.
@@
From the persistent and intermittent RFI masks, we produced a combined mask for each azimuth scan. We ultimately binned the data into 1 MHz bins for map-making (Section IV.3) and the subsequent Faraday synthesis. Any 83.3 kHz channels identified as containing persistent RFI were excluded from the binning, while channels with only intermittent RFI were included, with the contaminated integrations masked. With 12 83.3 kHz channels contributing to each 1 MHz bin, it was possible to retain useful data from these narrow channels even in cases where more than 50% of integrations were masked for intermittent RFI, provided the adjacent channels were not heavily contaminated. In some cases, a 1 MHz bin is on the edge of a wider band of persistent RFI, resulting in only a few channels contributing to the bin. In these cases, the 1 MHz channels suffer increased noise (Section V.3) and a few were discarded following visual inspection of the resulting maps (Section VI.1).
IV.2 Calibrating the Scans
@@
IV.2.1 Conversion to Noise-source Units
@@
Here $\text{PI}_{\text{noise}}$ is the ‘polarized intensity’ of the noise polynomial, $\text{PI}_{\text{noise}}=\sqrt{(\text{RL*}_{\text{noise}})^{2}+(\text{LR*}_{\text{noise}})^{2}}$ and $\tau_{\text{noise}}=\frac{1}{2}\tan^{-1}(\text{RL*}_{\text{noise}},\text{LR*}_{\text{noise}})$ is the corresponding phase. We chose the approach of converting the data into noise-source units in this manner rather than deriving a temperature-dependent gain correction because (i) the temperature measurements were taken at longer time intervals of once per 15 minutes compared to the 3-minute cadence of the noise source, and (ii) the temperature measurements were not recorded specifically along the RFoF connection. Following the conversion of the data into noise-source units, we masked the integrations that included the noise-source signal for subsequent processing steps. The Gaussian fits to the raster scans (Section III.2.2) were then done in noise-source units (Figure 10).
IV.2.2 Conversion to Jy beam-1
@@
We determined the noise-source-units-to-Jy beam-1 conversion factor as a function of frequency by comparing the Gaussian-fitted amplitudes of the raster scans in RR* and LL* to the flux density of the corresponding calibrator sources in Perley and Butler (2017).
@@
Two of the calibrators, Cas A and Tau A, are supernova remnants, and are known to change over time. We tested their suitability as calibrators using the Cyg A conversion factors to calibrate the Cas A and Tau A raster amplitudes for days when both Cyg A and one of these were observed. Cyg A is a bright radio galaxy and its flux is not expected to vary in time. From the resulting calibrated Cas A and Tau A spectra, we found that these sources have changed substantially since the observations for Perley and Butler (2017) and we rejected them for calibration purposes. When the fourth calibrator, the radio galaxy Vir A, was calibrated using the conversion factors developed for Cyg A, the resulting spectrum agreed with Perley and Butler (2017). However, the lower flux density of Vir A led to a noisy spectrum, and using conversion factors developed from Vir A raster scans would have introduced additional noise to the scans. Ultimately, we only used Cyg A raster scans to calibrate all azimuth scans.
@@
We produced survey-wide conversion factors by combining the fitted amplitudes for all Cyg A raster scans to produce a single set of RR* and LL* amplitude spectra. For the conversion spectra to be valid for all scans, the noise source was required to remain constant throughout the survey. Microwave noise sources are sensitive to temperature, and so the noise source was maintained at a controlled temperature above ambient (at $30\arcdeg$ C) in a well-insulated box. We tested the effectiveness of this in achieving the required stability by applying the master conversion factor spectra to all Cyg A and Vir A raster scans and finding the average ratio between the expected spectrum for the source from Perley and Butler (2017) and the measured source spectrum for each scan. Figure 14 shows the average ratio as a function of ambient temperature for each Cyg A (blue dots) and Vir A (red dots) scan. We note that although there is a slight correlation of this ratio with ambient temperature, the ratio remains within 3.5% of unity across the full range of temperatures. To account for the slight variation in the noise source across the survey, we applied a temperature-based scale factor to each scan. We determined the temperature scale factor from the linear fit to the average expected-to-measured ratio as a function of temperature values, shown as a solid black line in Figure 14.
@@
We applied the noise-source-units-to-Jy beam-1 conversion factors to the RR* and LL* components of all azimuth scans such that each of RR* and LL* represent the total intensity. After this calibration process the median percent difference between RR* and LL* for all scans was less than 5% across most of the frequency band, peaking at 10% near 450 MHz. We then calculated the Stokes $I$ total intensity values for the azimuth scans as $I=\frac{1}{2}(\text{RR*}+\text{LL*})$, following Wolleben et al. (2019, 2021). [^8] The correct definition of Stokes $I$ is $I=\text{RR*}+\text{LL*}$, without the factor of 1/2 (Robishaw and Heiles, 2021). However, our method of calibrating RR* and LL* intensities independently produces maps of total and polarized intensity that agree in brightness with previous radio surveys.
@@
At this stage, no further angle corrections were applied to RL* and LR*, and it was assumed that $Q=\text{RL*}$ and $U=\text{LR*}$ (similar to Wolleben et al., 2019, 2021). The noise-source-units-to-Jy beam-1 conversion factor, $R_{QU}$, used for Stokes $Q$ and $U$ is
@@
Figure 14: The ratio of the expected Perley and Butler (2017) flux density to the raster scan amplitudes calibrated using the survey-wide calibration factors averaged across the spectrum for each scan as a function of ambient temperature. Vir A scans are plotted with red squares and Cyg A scans are plotted with blue circles. A linear fit to the ratio versus temperature plot is shown as a black line.
@@
IV.3 Base level Adjustments and Filtering
Following RFI excision and calibration into Jy beam-1, we binned the Stokes $Q$, $U$ and $I$ data into 1 MHz frequency channels, each spanning 12 of the 83.3 kHz-wide channels. At this stage, we also separated the scans into east and west halves, which we then saved as SDFITS files to feed into the S-PASS software. Henceforth we use ‘scan’ to refer to an individual $180\arcdeg$ east or west scan rather than the full $360\arcdeg$ raw observation scans. We applied several initial steps to the full set of scans at each binned frequency to clean up the data prior to inclusion in the maps: